# Analyzing User Characteristics of Hate Speech Spreaders on Social Media

### v2

@@ 3.4.3. Effect estimation
We now model the relationships between the user attributes $A$ and the users’ probability to reshare hate speech $Y$ while controlling for PLV ${X}$. In our work, we are interested in understanding the effects of the following user attributes: verified ($=1$ if the user is verified, 00 otherwise), account age (in days), #posts (number of posts the user did), #followers, and #friends.
Explainable boosting machine: In our framework, we model the effects of user attributes $A$ on users’ probability to reshare hate speech $Y$, while controlling for PLV $X$, with an explainable boosting machine (EBM) (Nori et al., 2019). EBM is a tree-based, cyclic gradient boosting GAM (generalized additive model) with automatic interaction detection. It provides state-of-the-art performance while being fully explainable. Formally, we model
@@
3.5. Establishing causality
Our framework models the effect of user attributes on the probability to reshare hate speech while controlling for the past latent vulnerability (PLV) of users. Under certain assumptions, we can interpret our findings as causal effects. For this, we draw a causal graph in Figure 1, where the treatment is the user attribute and the outcome is the probability to reshare. Previous research has shown that the PLV influences both the user attributes and the probability to reshare (Benevenuto et al., 2009; Cheng et al., 2021), which then acts as the confounder.
Formally, we aim to identify the distribution of the potential outcome $Y_{u}({a})$ for each configuration of ${a}$ (for simplicity, we use $Y_{u}$ here, which can be replaced by $Y_{uc}$ for each hate speech cluster $c\in C$). However, this task is challenging due to the fundamental problem of causal inference (Pearl, 2015) in that we have only observed one potential outcome under one configuration of ${a}$ for each user (sometimes also referred to as observations being partially observable). Our task of estimating causal effects is based on $\mathbb{E}[Y_{u}({a})\,\mid\,{A}_{u}={a}]$. Note that this is a counterfactual estimate, which is different from a correlation estimate $\mathbb{E}[Y_{u}({a})]$, that is, $\mathbb{E}[Y_{u}({a})]\neq\mathbb{E}[Y_{u}({a})\,\mid\,{A}_{u}={a}]$. Because of that, we need to account for the confounder $X$, which simultaneously influences the treatment ${A}$ and the outcome $Y$. To make causal inferences, we use the PLV of users as the confounder $X$. When accounting for the confounder $X$, we can compute causal effects from observational data, i.e., $\mathbb{E}[Y_{u}\,\mid\,\mathrm{do}({a}),x]$.
Assumptions for causal inference: In addition, we make the following standard assumptions in causal inference (Rubin, 2005) to ensure identification:
  1. (1)
    Stable unit treatment value assumption (SUTVA): There is no interference between individuals and no different version of a cause. In the context of our work, SUTVA means that a user’s probability to reshare hate speech is independent of the attributes of other users (e.g., sharing hate speech does not depend on the follower count of other users). Moreover, it means that the values of the user attributes have the same interpretation across all users.
  2. (2)
    Positivity assumption: $10<p({a}_{u}\in A\,\mid\,x_{u})<$ 1, for all sets $A$ with $p(A)>0$. This ensures that all users could potentially receive treatment, given confounder $X$ (see Appendix E for additional analyses).
  3. (3)
    Ignorability: The confounder $X$ is sufficient to capture all variables that influence both treatments $A$ and outcome $Y$: $p({a}_{u}\in A\,\mid\,{X}_{u}(a_{1}),\ldots,{X}_{u}(a_{m}),x_{u})=p({a}_{u}\,\mid\,x_{u})$. This means the confounder $X$, here PLV, accounts for all confounding bias.
Importantly, the assumptions have been made earlier for modeling online user behavior (Cheng et al., 2021) and are reasonable given that our model uses the identical click data that users have for making resharing decisions.
Causal sensitivity analysis: To test for unobserved confounding, we perform a causal sensitivity analysis, a state-of-the-art tool (Frauen et al., 2023; Maarouf et al., 2024) that tests how strong an unobserved confounder would have to be to explain away the observed effects (see Appendix F for more details). Our analysis reveals that a potential unobserved confounder would have to be unreasonably strong to explain away our causal conclusions, which establishes that the estimated effects are causal.
Figure 1. Causal graph of our problem setting.
@@ 4.1. Data
Model variables: We analyze the effects of five user attributes on the probability to reshare: $\boldsymbol{a}_{1}$: verified ($=1$ if the user is verified, 0 otherwise); $\boldsymbol{a}_{2}$: account age (in days); $\boldsymbol{a}_{3}$: #post (number of posts by the user); $\boldsymbol{a}_{4}$: #follower (number of followers of the user); and $\boldsymbol{a}_{5}$: #friend (number of friends of the user). The distributions of #post, #follower, and #friend are highly right-skewed, because of which we log these user attributes. In the Appendix, Table 4 shows summary statistics for the different user attributes.
Ethical statement: Data collection was conducted by following standards for ethical research (Rivers and Lewis, 2014). Our analyses are based on publicly available data, and we only report aggregated results. We respect users’ privacy by not publishing usernames and do not attempt to track or de-identify users who decided to remain anonymous. We have received ethical approval by our ethics committee (=IRB) at the University of (anonymized). The approval number is EK-MIS-2023-160.
@@ 4.2. Baselines
In this work, we leverage a debiasing framework to model the effects of user attributes on users’ probability to reshare hate speech using an EBM. We refer to our framework as DF-EBM. Note that, for benchmarking, our setting requires baselines that provide explainable results since we aim to provide an understanding of the user attributes that make users reshare hate speech. To the best of our knowledge, we are aware of only one benchmark where the effects of user attributes on the probability to reshare content were modeled: (1) DF-naïve (Cheng et al., 2021) performs debiasing but, in contrast to our framework, is fully linear. Formally, instead of allowing for non-linear relationships in step  3, DF-naïve models linear effects of user attributes on the probability to reshare content. It is thus explainable as it outputs coefficients for each user attribute, but it cannot capture non-linearities and is thus not flexible since the modeled effects are assumed to be linear.
In addition, we design the following variants of our debiasing framework that we also use as baselines: (2) DF-ExNN where we replace the EBM with an explainable neural network (Yang et al., 2020). The explainable neural network enhances the explainability of a vanilla neural network by adding architecture constraints, which is supposed to achieve a balance between prediction performance and model explainability. (3) DF-NAM where we replace the EBM with a neural additive model (Agarwal et al., 2021), which learns a linear combination of neural networks that each attend to a single input feature. This way, it remains explainable while being able to model non-linear effects.
4.3. Performance metrics
For evaluation, we split our dataset into train/test set (80/20). Our performance metrics are two-fold. (1) To evaluate the performance of the entire framework (step  3), we use the root mean squared error (RMSE). (2) We additionally evaluate the performance of the recommendation via recall@$k$ and NDCG@$k$. The former is a measure of the posts that were interesting to the user out of those that the model predicted to be interesting as the top-$k$ posts. The latter measures the accuracy of the recommendation by comparing the ground-truth interestingness to the predicted ranking of posts for the top-$k$ ranked posts.
@@ 4.4. Implementation details
We implemented our causal framework in Python 3.8. For the hate speech clustering, we use BERTopic model all-MiniLM-L6-v2 (Grootendorst, 2022) to generate embeddings and UMAP to reduce the embeddings into a seven-dimensional space. To cluster, we use HDBSCAN with a minimum cluster size of 10. We name the resulting cluster with the LLAMA-3 model Llama-3-70b-chat-hf (Meta, 2024). To model the effects, we use the EBMRegressor of InterpretML (Nori et al., 2019). We allow for ten interactions in the EBM model and run it for a maximum of 5,000 rounds. Moreover, we performed hyperparameter tuning (see Appendix B). We average the performance of the EBMRegressor as well as our baselines over five runs. To model PLV, we use BPRMF (Rendle et al., 2009). We set the learning rate of our BPRMF model to 0.001, the batch size to 64, and the embedding size to 64. We average the performance over five runs with varying train/test splits. For the IPS reweighting with $\pi^{neural}_{h}$, we employ linear discriminant analysis (LDA) topic modeling [^2] https://scikit-learn.org/stable/modules/generated/sklearn.discriminant_analysis.LinearDiscriminantAnalysis.html to get a latent representation $e_{h}$ of the topics in the posts.
@@ 5.1. Hate speech cluster
We begin by reporting the hate speech clusters resulting from our topic modeling. We cluster 270 hate speech root posts into four clusters. The optimal number of clusters is determined automatically by HDBSCAN, which allows us more flexibility since we do not have to determine the number of clusters beforehand (Campello, Ricardo J. G. B. et al., 2013).
@@
5.2. Performance comparison
We next report the performance of our framework when predicting the effects of user attributes on the users’ probability to reshare hate speech. We compare our framework against multiple baselines based on the RMSE (root mean squared error). To the best of our knowledge, there is only one baseline in the literature: DF-naïve (Cheng et al., 2021), which builds upon a linear approach and thus can not account for non-linearites. All of the baselines as well as our model are explainable since we aim to provide an understanding of the user attributes that make users reshare hate speech.
Table 1 shows the RMSE for each variant of our framework. Our debiasing framework DF-EBM outperforms the DF-naïve baseline with a relative improvement of 2.5%. The improvement is statistically significant according to Welch’s $t$-test (Welch, 1947) at a $p<0.001$ level. Importantly, both frameworks have access to identical data so the performance gain must be attributed solely to the flexibility of our framework in capturing nonlinearities. Moreover, our DF-EBM also outperforms the other custom baselines that are non-linear. Here, the Welch’s $t$-tests (Welch, 1947) are again statistically significant at the $p<0.01$ level. Together, this demonstrates that our debasing framework DF-EBM is superior.
@@ 5.2. Performance comparison
DF-naïve [Uncaptioned image] [Uncaptioned image] 0.0885
@@
Table 1. Performance comparison (measured by the RMSE) for predicting users’ probability to reshare hate speech. Our debiasing framework performs best (relative improvement of 2.5% over the best baseline, whereby DF-naïve is the only baseline that we are aware of in the literature).
@@ 5.3. Analyzing hate speech resharing
We now leverage the explainability of our framework to understand which user attributes affect the users’ probability to reshare hate speech. For this, we analyzed the effects of users’ verified status, account age, number of posts, number of followers, and number of friends. We first look at the feature importance for each user attribute, which is computed as the average absolute predicted value of each feature for the training set, i.e. for each feature, we predict the absolute value for each sample in the training set and then average these scores. This is weighted by the density of the train set. Here, we find that the number of posts is the most influential user attribute when predicting users’ probability to reshare hate speech (see Figure 2). Moreover, the second most influential attribute is the number of followers. Thereby, we add to the literature, where the popularity of a user and their activity level are important indicators as to whether they post hate speech (Hua et al., 2020; Ribeiro et al., 2018), whereas we show that the number of posts and followers also cause a larger propensity to reshare hate speech.
We next look at the effects of each user attribute on the probability to reshare hate speech. Figure 3 shows the contribution functions for the user attributes. The contribution is shown on the same scale as the output variable and the graph is normalized so that the average prediction is at 0. For example, a score of $-0.1$ at an account age of 100 days would mean that a user with a 100-day-old account is 10 percentage points less likely to reshare hate speech than the average prediction. Overall, we find that users with fewer posts, fewer followers, and fewer friends are more likely to reshare hate speech. In return, newer accounts are less likely to do so. Moreover, we find that verified users are 0.016 (±$\pm 0.004$) percentage points more likely to reshare hate speech than the average prediction, which is contrary to our expectations that a verified status is typically linked to users with a large reputation who should thus engage less with hate speech.
@@ 5.5. Ablation study
5.6. Robustness checks
Modeling past latent vulnerability: We perform a robustness check to validate our choice of how we model the past latent vulnerability (PLV). Recall that we model PLV using the recommendation algorithm BPRMF. However, we have multiple ways to debias propensity scores, implying that we can use multiple inputs to the BPRMF model. Here, we check the prediction performance of the BPRMF with different propensity score inputs. We measure the prediction performance using recall@$k$ and NDCG@$k$. The former measures how many of the predicted posts were interesting to the user and the latter measures the accuracy of the prediction. This way, we can get an estimate of how well the recommendation algorithm was able to model PLV for each user. Table 3 shows the prediction performance of BPRMF with different inputs. We find that the BPRMF algorithms with debiased propensity scores are better at predicting hate speech that the user is vulnerable to. This confirms our choice to employ IPS reweighting in our debiasing framework when modeling PLV. We further perform a comparison of PLV to hate speech and PLV to normal content (see Appendix D for the complete analysis).
Evaluation with semi-synthetic data: In causal inference, the performance of effect estimation should be evaluated on counterfactuals. Since these are missing in observational data, i.e. the fundamental problem of causal inference, we follow common practice (Frauen et al., 2023) and generate a semi-synthetic dataset (see Appendix G for more details). We find that our framework DF-EBM still outperforms baselines and hence confirm the effectiveness of our debiasing framework.
recall@$k$
Model $k=20$ $k=40$ $k=60$ $k=80$
BPRMF 19.35 30.90 36.20 40.17
BPRMF-V 18.02 31.18 36.76 40.69
BPRMF-F 19.48 32.11* 37.92* 42.16*
BPRMF-NN 19.23 31.16 36.62 40.80
(a) Performance evaluated based on recall@$k$
NDCG@$k$
Model $k=20$ $k=40$ $k=60$ $k=80$
BPRMF 7.02 9.39 10.33 10.97
BPRMF-V 6.67 9.37 10.36 11.00
BPRMF-F 7.04 9.63 10.66 11.35*
BPRMF-NN 6.82 9.27 10.24 10.92
(b) Performance evaluated based on NDCG@$k$
@@ 6. Discussion
When looking at the different types of hate speech, we make the interesting observation that other dynamics spread racial and misogynistic hate than political hate. Users with little social influence tend to spread more racial and misogynistic hate while users with high social influence tend to spread political hate. This might be the case because racial and misogynistic hate is socially unacceptable and circulates primarily in small online communities where extreme views are normalized and users do not have to fear backlash for resharing. In return, the political landscape is widely discussed online and political hate might border what is still deemed acceptable to express online or is harder to detect for moderation efforts.
@@
Implications: Our framework shows the influences of user attributes on the probability to reshare hate speech and therefore generates insights on how to take action against the resharing of hate speech. For example, online platforms might consider not showing the popularity of a user by removing follower counts from the profile information. This way, users are unaware of how many people see their posts, which might lead them to be more selective in the content they share. Platforms could also vary their moderation measures depending on the user’s follower count, e.g., decrease the response times to reports against hate speech from popular users. However, this would require a shift in resources and could raise ethical considerations about the uneven moderation of users. Given the positive effects of a verified status on hate speech resharing, online platforms might also consider removing the verified status. This might lead to users being more responsible in the content they reshare.
Conclusion: In this work, we propose a novel framework to model the effects of user attributes on the probability to reshare hate speech while controlling for the past latent vulnerability to hate speech. Our framework provides flexible and explainable results while achieving state-of-the-art performance. We also cluster hate speech into its multiple types, which allows us to model the resharing dynamics individually.
Acknowledgements.
Funding by the German Research Foundation (Grant: 543018872) is acknowledged.
@@ Appendix B Hyperparameter tuning
Appendix C Robustness check for smoothing operator
In IPS for reweighting observational social media data, the IPS for rare events could be very high. A smoothing operator can lead to more stable propensity scores and prevent overfitting. In this work, $\mu$ is the smoothing operator. We chose $\mu=0.5$ since popularity-related measures follow a power law distribution, i.e., $\mu\leq 1$. We performed a new robustness check to study the sensitivity of the recommendation algorithm with $\mu=[0.1,0.5,1.0]$ for the virality and the follower-based propensity score. The results with respect to recall@$k$ are in Table 6. We found that the overall findings remain robust.

### v3

@@ 3.4.2. Effect estimation
We now model the relationships between the user attributes $A$ and the users’ probability to reshare hate speech $Y$ while controlling for PLV ${X}$ for each user $u$. We model the effects using an explainable boosting machine (EBM) (Nori et al., 2019). EBM is a tree-based, cyclic gradient boosting GAM (generalized additive model) with automatic interaction detection. It provides state-of-the-art performance while being fully explainable.
In our work, we are interested in understanding the effects of the following user attributes $a\in A$: verified ($=1$ if the user is verified, 00 otherwise), account age (in days), #posts (number of posts the user did), #followers, and #friends. Formally, we model
@@ 4.1. Data
Model variables: We analyze the effects of five user attributes on the probability to reshare: $\boldsymbol{a}_{1}$: verified ($=1$ if the user is verified, 0 otherwise); $\boldsymbol{a}_{2}$: account age (in days); $\boldsymbol{a}_{3}$: #post (number of posts by the user); $\boldsymbol{a}_{4}$: #follower (number of followers of the user); and $\boldsymbol{a}_{5}$: #friend (number of friends of the user). The distributions of #post, #follower, and #friend are highly right-skewed, because of which we log these user attributes. Appendix Table 3 shows summary statistics for the different user attributes.
Ethical statement: Data collection was conducted by following standards for ethical research (Rivers and Lewis, 2014). Our analyses are based on publicly available data, and we only report aggregated results. We respect users’ privacy by not publishing usernames and do not attempt to track or de-identify users who decided to remain anonymous. We have received ethical approval by our ethics committee (=IRB) at the University of (anonymized). The approval number is (anonymized).
@@ 4.2. Baselines
In this work, we leverage a debiasing framework to model the effects of user attributes on users’ probability to reshare hate speech using an EBM. We refer to our framework as DF-EBM. Note that, for benchmarking, our setting requires baselines that provide explainable results since we aim to provide an understanding of the user attributes that make users reshare hate speech. To the best of our knowledge, we are aware of only one benchmark where the effects of user attributes on the probability to reshare content were modeled: (1) DF-linear (Cheng et al., 2021) performs debiasing but, in contrast to our framework, is fully linear. Formally, instead of allowing for non-linear relationships in step  3, DF-linear models linear effects of user attributes on the probability to reshare content. It is thus explainable as it outputs coefficients for each user attribute, but it cannot capture non-linearities and is thus not flexible since the modeled effects are assumed to be linear.
In addition, we design the following variants of our debiasing framework that we also use as baselines: (2) DF-ExNN where we replace the EBM with an explainable neural network (Yang et al., 2020). The explainable neural network enhances the explainability of a vanilla neural network by adding architecture constraints, which is to achieve a balance between prediction performance and model explainability. (3) DF-NAM where we replace the EBM with a neural additive model (Agarwal et al., 2021), which learns a linear combination of neural networks that each attend to a single input feature. This way, it remains explainable while being able to model non-linear effects.
@@ 4.3. Implementation details
We implemented our framework in Python 3.8. For the hate speech clustering, we use BERTopic model all-MiniLM-L6-v2 (Grootendorst, 2022) to generate embeddings and UMAP to reduce the embeddings into a seven-dimensional space. To cluster, we use HDBSCAN with a minimum cluster size of 10. We name the resulting cluster with the LLAMA-3 model Llama-3-70b-chat-hf (Meta, 2024). To model the effects, we use the EBMRegressor of InterpretML (Nori et al., 2019). We allow for ten interactions in the EBM model and run it for a maximum of 5,000 rounds. Moreover, we performed hyperparameter tuning (see Appendix B). For evaluation, we split our dataset into train/test set (80/20). To evaluate the performance of the entire framework (step  3), we use the root mean squared error (RMSE). We average the performance of the EBMRegressor as well as our baselines over five runs. To model PLV, we use BPRMF (Rendle et al., 2009). We set the learning rate of our BPRMF model to 0.001, the batch size to 64, and the embedding size to 64. We average the performance over five runs with varying train/test splits. For the IPS reweighting with $\pi^{neural}_{h}$, we employ linear discriminant analysis (LDA) topic modeling [^2] https://scikit-learn.org/stable/modules/generated/sklearn.discriminant_analysis.LinearDiscriminantAnalysis.html to get a latent representation $e_{h}$ of the topics in the posts.
@@ 5.1. Hate speech cluster
We begin by reporting the hate speech clusters resulting from our topic modeling. We cluster the 270 hate speech root posts into four clusters, which we labeled automatically. The optimal number of clusters is determined automatically by HDBSCAN, which allows us more flexibility since we do not have to determine the number of clusters beforehand (Campello, Ricardo J. G. B. et al., 2013).
@@
DF-linear [Uncaptioned image] [Uncaptioned image] 0.0885
@@
Table 1. Performance comparison (measured by the RMSE) for predicting users’ probability to reshare hate speech. Our debiasing framework performs best (relative improvement of 2.5% over the best baseline, whereby DF-linear is the only baseline that we are aware of in the literature).
5.2. Performance comparison
We next report the performance of our framework when predicting the effects of user attributes on the users’ probability to reshare hate speech. We compare our framework against multiple baselines based on the RMSE (root mean squared error). To the best of our knowledge, there is only one baseline in the literature: DF-linear (Cheng et al., 2021), which builds upon a linear approach and thus can not account for non-linearites. All of the baselines as well as our model are explainable since we aim to provide an understanding of the user attributes that make users reshare hate speech.
Table 1 shows the RMSE for each variant of our framework. Our debiasing framework DF-EBM outperforms the DF-linear baseline with a relative improvement of 2.5%. The improvement is statistically significant according to Welch’s $t$-test (Welch, 1947) at a $p<0.001$ level. Importantly, both frameworks have access to identical data so the performance gain must be attributed solely to the flexibility of our framework in capturing nonlinearities. Moreover, our DF-EBM also outperforms the other custom baselines that are non-linear. Here, the Welch’s $t$-tests (Welch, 1947) are again statistically significant at the $p<0.01$ level. Together, this demonstrates that our debasing framework DF-EBM is superior.
@@ 5.3. Analyzing hate speech resharing
We now leverage the explainability of our framework to understand which user attributes affect the users’ probability to reshare hate speech. For this, we analyzed the effects of users’ verified status, account age, number of posts, number of followers, and number of friends. We first look at the feature importance for each user attribute, which is computed as the average absolute predicted value of each feature for the training set (i.e., for each feature, we predict the absolute value for each sample in the training set and then average these scores). This is weighted by the relative frequency of each feature in the train set. Here, we find that the number of posts is the most influential user attribute when predicting users’ probability to reshare hate speech (see Figure 1). Moreover, the second most influential attribute is the number of followers. Thereby, we add to the literature where the popularity of a user and their activity level are important indicators as to whether they post hate speech (Hua et al., 2020; Ribeiro et al., 2018), whereas we show that the number of posts and followers also cause a larger propensity to reshare hate speech.
We next look at the effects of each user attribute on the probability to reshare hate speech. Figure 2 shows the contribution of each user attribute. The contribution is shown on the same scale as the output variable, and the graph is normalized so that the average prediction is at 0. For example, a score of $-0.1$ at an account age of 100 days would mean that a user with a 100-day-old account is 10 percentage points less likely to reshare hate speech than the average prediction. Overall, we find that users with fewer posts, fewer followers, and fewer friends are more likely to reshare hate speech. In return, newer accounts are less likely to do so. Moreover, we find that verified users are 0.016 (±$\pm 0.004$) percentage points more likely to reshare hate speech than the average prediction, which is contrary to our expectations that a verified status is typically linked to users with a large reputation who should thus engage less with hate speech. We further perform a comparison of PLV to hate speech and PLV to normal content (see Appendix E for the complete analysis).
@@ 6. Discussion
When looking at the different types of hate speech, we make the interesting observation that the dynamics behind the of spread racial and misogynistic hate and political hate differ. Users with little social influence tend to spread more racial and misogynistic hate while users with high social influence tend to spread political hate. This might be the case because racial and misogynistic hate is socially unacceptable and circulates primarily in small online communities where extreme views are normalized and users do not have to fear backlash for resharing. In return, the political landscape is widely discussed online and political hate might border what is still deemed acceptable to express online or is harder to detect for moderation efforts.
@@
Implications: Our framework shows the influences of user attributes on the probability to reshare hate speech and therefore generates insights on how to take action against the resharing of hate speech. For example, online platforms might consider not showing the popularity of a user by removing follower counts from the profile information. This way, users are unaware of how many people see their posts, which might lead them to be more selective in the content they share.
Platforms could also vary their moderation measures depending on the user’s follower count, e.g., decrease the response times to reports against hate speech from popular users. However, this would require a shift in resources and could raise ethical considerations about the uneven moderation of users. Given the positive effects of a verified status on hate speech resharing, online platforms might also consider removing the verified status. This might lead to users being more responsible in the content they reshare.
Conclusion: In this work, we leverage a state-of-the-art debiasing framework to model the effects of user attributes on the probability to reshare hate speech while controlling for the past latent vulnerability to hate speech. Our framework provides flexible and explainable results while achieving state-of-the-art performance. We also cluster hate speech into its multiple types, which allows us to model the resharing dynamics individually.
@@ Appendix B Hyperparameter tuning
Appendix C Insights into the different types of hate speech
[…]
Hate speech cluster Reshares Characteristic words
Racist and misogynistic rants 15.219 hate, niggas, nigga, dont, yall, like, ass, ugly, bitch, fucking
Anti-Trump and anti-right wing rants 6.207 idiot, trump, isis, idiots, syria, says, attack, obama, war, run
Toxic rants 512 dont, women, fucking, brutal, disgusting, life, white, just, bad, lifequotes
Social media noise 51 summer, tomorrow, selby, model, focus, views, hes, make, time, just
Table 5. Number of reshares and characteristic words of each hate speech cluster defined as the top-10 words with regard to c-TF-IDF weighting.
Hate speech cluster Example posts
[…]
Table 6. Example posts from each hate speech cluster
Appendix D Robustness checks
Modeling past latent vulnerability: We perform a robustness check to validate our choice of how we model the past latent vulnerability (PLV). Recall that we model PLV using the recommendation algorithm BPRMF. However, we have multiple ways to debias propensity scores, implying that we can use multiple inputs to the BPRMF model. Here, we check the prediction performance of the BPRMF with different propensity score inputs. We measure the prediction performance using recall@$k$ and NDCG@$k$. The former measures how many of the predicted posts were interesting to the user out of those that the model predicted to be interesting as the top-$k$ posts. The latter measures the accuracy of the prediction by comparing the ground-truth interestingness to the predicted ranking of posts for the top-$k$ ranked posts. This way, we can get an estimate of how well the recommendation algorithm was able to model PLV for each user.
Table 7 shows the prediction performance of BPRMF with different inputs. We find that the BPRMF algorithms with debiased propensity scores are better at predicting hate speech that the user is vulnerable to. This confirms our choice to employ IPS reweighting in our debiasing framework when modeling PLV.
recall@$k$
Model $k=20$ $k=40$ $k=60$ $k=80$
BPRMF 19.35 30.90 36.20 40.17
BPRMF-V 18.02 31.18 36.76 40.69
BPRMF-F 19.48 32.11* 37.92* 42.16*
BPRMF-NN 19.23 31.16 36.62 40.80
(a) Performance evaluated based on recall@$k$
NDCG@$k$
Model $k=20$ $k=40$ $k=60$ $k=80$
BPRMF 7.02 9.39 10.33 10.97
BPRMF-V 6.67 9.37 10.36 11.00
BPRMF-F 7.04 9.63 10.66 11.35*
BPRMF-NN 6.82 9.27 10.24 10.92
(b) Performance evaluated based on NDCG@$k$
Robustness check for smoothing operator In IPS for reweighting observational social media data, the IPS for rare events could be very high. A smoothing operator can lead to more stable propensity scores and prevent overfitting. In this work, $\mu$ is the smoothing operator. We chose $\mu=0.5$ since popularity-related measures follow a power law distribution, i.e., $\mu\leq 1$. We performed a new robustness check to study the sensitivity of the recommendation algorithm with $\mu=[0.1,0.5,1.0]$ for the virality and the follower-based propensity score. The results with respect to recall@$k$ are in Table 8. We found that the overall findings remain robust.